A diagonal walk through a small rectangular garden 9 meters by 12 meters can be built at $10 per linear meter. How much will the walk cost? ...
Answer by ewatrrr(24785)


Hi,
rectangular garden 9 meters by 12 meters
Finding the length of the diagonal d
Applying the Pythagorean Theorem
9^2 + 12^2 = d^2
225 = d^2
+-15 = d (discarding the negative solution)
15m = d, the length of the diagonal walk
How much will the walk cost at $10/linear meter
15m*$10 = $150